Need help with below:

Twenty metrics of liquidity, Solvency, and Profitability

The comparative financial statements of Automotive Solutions Inc. are as follows. The market price of Automotive Solutions Inc. common stock was $119.70 on December 31, 20Y8.

AUTOMOTIVE SOLUTIONS INC.
Comparative Income Statement
For the Years Ended December 31, 20Y8 and 20Y7
    20Y8     20Y7
Sales $10,000,000 $ 9,400,000
Cost of goods sold (5,350,000) (4,950,000)
Gross profit $ 4,650,000 $ 4,450,000
Selling expenses $ (2,000,000) $(1,880,000)
Administrative expenses (1,500,000) (1,410,000)
Total operating expenses $ (3,500,000) $(3,290,000)
Operating income $ 1,150,000 $ 1,160,000
Other revenue and expense:
    Other revenue 150,000 140,000
    Other expense (interest) (170,000) (150,000)
Income before income tax $ 1,130,000 $ 1,150,000
Income tax expense (230,000) (225,000)
Net income $ 900,000 $ 925,000
AUTOMOTIVE SOLUTIONS INC.
Comparative Statement of Stockholders’ Equity
For the Years Ended December 31, 20Y8 and 20Y7
20Y8 20Y7
Preferred
Stock
Common
Stock
Retained
Earnings
Preferred
Stock
Common
Stock
Retained
Earnings
Balances, Jan. 1 $500,000 $500,000 $5,375,000 $500,000 $500,000 $4,545,000
Net income 900,000 925,000
Dividends:
    Preferred stock (45,000) (45,000)
    Common stock (50,000) (50,000)
Balances, Dec. 31 $500,000 $500,000 $6,180,000 $500,000 $500,000 $5,375,000
AUTOMOTIVE SOLUTIONS INC.
Comparative Balance Sheet
December 31, 20Y8 and 20Y7
    Dec. 31, 20Y8     Dec. 31, 20Y7
Assets
Current assets:
    Cash $ 500,000 $ 400,000
    Marketable securities 1,010,000 1,000,000
    Accounts receivable (net) 740,000 510,000
    Inventories 1,190,000 950,000
    Prepaid expenses 250,000 229,000
        Total current assets $3,690,000 $3,089,000
Long-term investments 2,350,000 2,300,000
Property, plant, and equipment (net) 3,740,000 3,366,000
Total assets $9,780,000 $8,755,000
Liabilities
Current liabilities $ 900,000 $ 880,000
Long-term liabilities:
    Mortgage note payable, 10% $ 200,000 $ 0
    Bonds payable, 10% 1,500,000 1,500,000
        Total long-term liabilities $1,700,000 $1,500,000
Total liabilities $2,600,000 $2,380,000
Stockholders' Equity
Preferred $0.90 stock, $10 par $ 500,000 $ 500,000
Common stock, $5 par 500,000 500,000
Retained earnings 6,180,000 5,375,500
Total stockholders' equity $7,180,000 $6,375,000
Total liabilities and stockholders' equity $9,780,000 $8,755,000

Instructions:

Determine the following measures for 20Y8.

Round ratio values to one decimal place and dollar amounts to the nearest cent. For number of days' sales in receivables and number of days' sales in inventory, round intermediate calculations to the nearest whole dollar and final amounts to one decimal place. Assume there are 365 days in the year.

1. Working capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Days' sales in receivables days
6. Inventory turnover
7. Days' sales in inventory days
8. Debt ratio %
9. Ratio of liabilities to stockholders' equity
10. Ratio of fixed assets to long-term liabilities
11. Times interest earned times
12. Times preferred dividends earned times
13. Asset turnover
14. Return on total assets %
15. Return on stockholders’ equity %
16. Return on common stockholders’ equity %
17. Earnings per share on common stock $
18. Price-earnings ratio
19. Dividends per share of common stock $
20. Dividend yield %
Answer #1

Answers:

1. Working Capital $      2,790,000.00
2. Current ratio                         4.1
3. Quick ratio                         2.5
4. Accounts receivable turnover                       16.0
5. Days sales in receivables           22.8 days
6. Inventory turnover                         5.0
7. Days sales in inventory           73.0 days
8. Debt ratio           26.6 %
9. Ratio of liabilities to stockholders equity                         0.4
OR           36.2 %
10. Ratio of fixed assets to long term liabilities                         2.2
OR
        220.0 %
11. Times interest earned             6.8 times
12. Times preferred dividends earned           20.0 times
13. Asset turnover                         1.1
14. Return on total assets             9.7 %
15. Return on stockholder's equity           13.3 %
16. Return on common stockholder's equity           13.6 %
17. Earnings per share on common stock $                       8.55
18. Price earnings ratio                       14.0
19. Dividends per share of common stock $                       0.50
20. Dividend yield             0.4 %

Calculations and formulas:

1. Working Capital $ =3690000-900000
2. Current ratio =3690000/900000
3. Quick ratio =(3690000-1190000-250000)/900000
4. Accounts receivable turnover =10000000/((740000+510000)/2)
5. Days sales in receivables =365/16 days
6. Inventory turnover =5350000/((1190000+950000)/2)
7. Days sales in inventory =365/5 days
8. Debt ratio =2600000/9780000*100 %
9. Ratio of liabilities to stockholders equity =2600000/7180000
OR =J27*100 %
10. Ratio of fixed assets to long term liabilities =3740000/1700000
OR
=J29*100 %
11. Times interest earned =1150000/170000 times
12. Times preferred dividends earned =900000/45000 times
13. Asset turnover =10000000/((9780000+8755000)/2)
14. Return on total assets =900000/((9780000+8755000)/2)*100 %
15. Return on stockholder's equity =900000/((7180000+6375000)/2)*100 %
16. Return on common stockholder's equity =(900000-45000)/((7180000+6375000-500000-500000)/2)*100 %
17. Earnings per share on common stock $ =(900000-45000)/((100000+100000)/2)
18. Price earnings ratio =119.7/J38
19. Dividends per share of common stock $ =50000/100000
20. Dividend yield =J40/119.7*100 %